OpenStudy (anonymous):

solve the system by the elimination method 5x+3y=-7 7x-2y=17

OpenStudy (anonymous):

We need to get rid of one of the variables. Since the y's have both a + and a -, let's get rid of them. Multiply the top equation by 2 and the bottom equation by 3 10x + 6y = -14 21x - 6y = 51 Now add the two equations, notice the y's drop out 31x = 37 divide by 31 x = 37/31 Now to find y, chose one of the two original equations I'll use the first one, it doesn't matter because this is not a "nice" answer. 5(37/31) + 3y = -7 185/31 + 3y = -7 subtract 185/31 from both sides 3y = -7 - 185/31 (We need common denominators) 3y = -217/31 - 185/31 3y = -402/31 divide by 3 (or multiply by 1/3) y = -134/31 So the point of intersection is (37/31 , -134/31) We should check in the other equation but since these numbers are so miserable, I used a calculator (using matrices, but that is for another time) Hope it helps.

OpenStudy (anonymous):

5x+3y=-7 mult this by 2 7x-2y=17 mult this by 3 gives 10x+6y=-14 21x-6y=51 add them we get ------------ 31x +0 =37 x= 37/31 ans subs this to the ist eq, 5x+3y=-7 5(37/31)+3y=-7 165/31 +3y=-7 3y= -7 - 165/31 y=402/(31x3) =-134/31 ans (37/31, -134/31) ans